# Emergent Hydrodynamics in an Exclusion Process with Long-Range Interactions

### v1

@@ 5 Conclusion
Several open directions now suggest themselves. First, uncovering a genuinely “non-local Kardar-Parisi-Zhang (KPZ)” regime would be particularly exciting. Introducing a weak asymmetry or external noise is a natural route, yet it remains unclear which—if any—perturbations actually generate KPZ-type scaling. A link between free fermions to which the SDEP maps and the KPZ universality class has been pointed out recently [68]. At the PDE level, a Hilbert-kernel variant of the KPZ equation has already been proposed earlier [69]; clarifying its connection to the SDEP is an open problem. Second, applying macroscopic-fluctuation theory, as recently done for the continuum Dyson [51] gas, should reveal how the hard-core constraint alters current large deviations; the long-range MFT developed in [70] offers a natural starting point. Third, coupling the SDEP to particle reservoirs could help us explore basic questions such as how the boundary-induced phase transitions that are well understood for short range interactions [71] would become in the presence of long range interacting. A few promising attempts are already present in the literature [72, 73]. Fourth, a full exploitation of the XX-chain mapping promises refined analytic control. This direction is currently being developed in ongoing work. Finally, conditioning the SSEP on finite (rather than maximal) activity, which corresponds to a Doob transform of the XXZ generator, remains a challenging but exciting avenue.
@@ Appendix A Conditioning a continuous-time Markov process on large deviations
Let $O\subset\{(i,j)\in\Omega\times\Omega:i\neq j\}$ be a specified set of ordered pairs of states (interpreted as jumps $i\to j$). Define the (extensive) counting process
@@
has a unique solution $s=s(j)$, and the long-time path measure of the original process conditioned on $J_{t}/t\approx j$ (microcanonical) is logarithmically equivalent to the path measure of the Doob-transformed process $M^{\mathrm{cond}}(s(j))$ (canonical). That is, bulk observables computed in either ensemble agree in the $t\to\infty$ limit; see [17, 18] and [75] for details.

### v2

@@ 5 Conclusion
Several open directions now suggest themselves. First, uncovering a genuinely “non-local Kardar-Parisi-Zhang (KPZ)” regime would be particularly exciting. Introducing a weak asymmetry or external noise is a natural route, yet it remains unclear which—if any—perturbations actually generate KPZ-type scaling. A link between free fermions to which the SDEP maps and the KPZ universality class has been pointed out recently [39]. At the PDE level, a Hilbert-kernel variant of the KPZ equation has already been proposed earlier [44]; clarifying its connection to the SDEP is an open problem. Second, applying macroscopic-fluctuation theory, as recently done for the continuum Dyson [24] gas, should reveal how the hard-core constraint alters current large deviations; the long-range MFT developed in [11] offers a natural starting point. Third, coupling the SDEP to particle reservoirs could help us explore basic questions such as how the boundary-induced phase transitions that are well understood for short range interactions [37] would change in the presence of long range interactions. A few promising attempts are already present in the literature [14, 60]. Fourth, a full exploitation of the XX-chain mapping promises refined analytic control. This direction is currently being developed in ongoing work. Finally, conditioning the SSEP on finite (rather than maximal) activity, which corresponds to a Doob transform of the XXZ generator, remains a challenging but exciting avenue.
@@ Appendix A Conditioning a continuous-time Markov process on large deviations
Let $O\subset\{(i,j)\in\Omega\times\Omega:i\neq j\}²$ be a specified set of ordered pairs of states (interpreted as jumps $i\to j$). Define the (extensive) counting process
@@
has a unique solution $s=s(j)$, and the long-time path measure of the original process conditioned on $J_{t}/t\approx j$ (microcanonical) is logarithmically equivalent to the path measure of the Doob-transformed process $M^{\mathrm{cond}}(s(j))$ (canonical). That is, bulk observables computed in either ensemble agree in the $t\to\infty$ limit; see [40, 17] and [59, 18] for details.